# EdgeAIGuard: Agentic LLMs for Minor Protection in Digital Spaces

### v1

@@ I Introduction
  • A Novel multi-agent framework EdgeAIGuard for minor’s online safety.
@@ II Related Work
The LLM-based approaches [33, 17, 12, 34, 11, 13] surpassing CNN and RNN models in detecting hate speech content in textual data. Authors in [33] explored the automatic detection of white supremacist hate speech on Twitter using the BERT architecture, which outperformed BiLSTM. Similarly, [17] introduced HateTinyLLM for efficient hate speech detection, using fine-tuned, decoder-only tiny large language models (tinyLLMs) combined with Low-Rank Adaptation (LoRA) and adapter methods. In [12], the authors presented SHIELD, an LLM-based technique that extracts rationales from input text to improve the interpretability of hate speech detection models. In [34], the researchers focused on enhancing hate speech detection with multilingual corpora and Cross-Lingual Learning (CLL) techniques, specifically for political discourse. Similarly, [11] investigated LLM-based methods for detecting sexist and hateful online content through information retrieval, employing zero-shot and few-shot learning approaches. Finally, [13] introduced TurkishBERTweet, a LLM-based method trained on over 894 million Turkish tweets.
Current research methods on hate speech detection using machine learning, deep learning, and LLM-based methods have significant limitations while deploying on edge devices. Traditional models such as SVM [22] and feature-based methods [26, 28] struggle with context-dependent hate speech and require extensive feature engineering. Deep learning models like CNNs, RNNs, and LSTMs [31, 30, 8] improve classification compared to traditional approaches but suffer from high computational costs and limited generalization beyond their training datasets. While LLM-based methods [33, 11, 17] show promising results, they have significant latency issues, making real-time detection impractical. Due to high computational demands, existing LLM-based approaches are inefficient for edge deployment, raising concerns about privacy, cost, and real-time inference. Additionally, existing methods rely on specific datasets and fail to adapt to evolving online threats, limiting real-world applicability. We propose an Agentic AI-based approach designed on an edge-first architecture to address these challenges, ensuring enhanced privacy, minimal latency, and adaptive threat detection.
@@ III Methodology
The EdgeAIGuard framework comprises four essential layers, which are designed to optimize the protection pipeline for minors. The input layer collects data and processes information from various digital platforms. The form of information can be extended to the ones shown in Figure 2, such as comments on a blog, discord chat, or any other platform that can be used for communication with minors. The edge processing unit forms the analytical core of the EdgeAIGuard framework, housing both the agent network and LLM engine. Context-aware decision-making is performed through local storage, which maintains emerging patterns and historical data for the AI agents. Finally, the intervention in the case of threat detection is executed by the protection layer in synchronization with one of the AI agents.
The proposed system adopts a sequential-parallel hybrid processing approach by allowing multiple specialized agents to process incoming data and information from the communications simultaneously. The hybrid approach maintains critical sequential dependencies essential for decision-making. It also optimizes the accuracy and processing speed, thus making the proposed EdgeAIGuard compliant with the rapid response time requirement for threat detection systems.
III-A Preliminaries and Problem Formulation
@@ III-B Input Layer
III-C Edge Processing Unit
@@
The notation $H$ represents the historical context vector, $g$ is the scoring function, and the operator $\circ$ represents the fusion operation. Lastly, we define an intervention agent that determines the most appropriate response strategy based on the sentinel and context agent inputs, respectively. We formally define the intervention agent in Equation 7.
@@ III-D Local Storage
The term $HC$ refers to the history cache, which utilizes a decay factor $\alpha$ and the current state represented by $CS(t)$, whereas the $pm_{z}$ refers to the set of patterns such that $(pattern,frequency,severity)$. The history cache maintains a record of recent system and communication states and is adjusted through time-weightage. The history cache also utilizes $\alpha$ to gradually reduce the influence of older information. However, it preserves important patterns accordingly. On the other hand, pattern memory stores the threat patterns that are recurring in nature, along with their severity and frequency. The memory enables the system to recognize evolving threat patterns and to adapt the responses.
III-E Protection Layer
@@
The notations $TD$ refers to the threat decision, $risk$ corresponds to the historical risks, and $\delta$ refers to the decision function. The weights for each of the terms are optimized so that the decision process for intervention is optimal. Subsequently, the intervention selection process is designed to balance the effectiveness of potential interventions while analyzing the impact on the user experience. In this regard, the intervention process employs a utility-based approach, as shown in Equation 13.
@@
III-F Objective Functions and Optimization
@@
where $loss_{ce}$ refers to the computational efficiency loss and $\lambda$s are the corresponding weights. We define each loss component in Equations 15, 16, 17, and 18 as follows:
@@ IV-A Experimental Setup
The EdgeAIGuard was evaluated using a combination of software simulations, controlled testbeds, and benchmark datasets to replicate real-world scenarios. We used NVIDIA Jetson AGX Orin to emulate the edge device, representing real-world deployment conditions due to its balance between computational power and energy efficiency. We also used Intel Xeon Gold 6226R CPU paired with an NVIDIA A100 GPU, representing a high-performance cloud environment for baseline comparisons. As mentioned in the algorithm and architectural workflow, we would leverage DeepSeek-7B (distilled version) [35], which is optimized for edge deployment without compromising performance. For the model development and fine-tuning, we employed PyTorch [36] and HuggingFace Transformers [37].
@@ IV-B Dataset Details
We mainly used two datasets to train the sentinel and context agents to identify implicit grooming attempts and detect harmful interactions. We utilized the online grooming dataset [41] that contains 1.2 million chat messages from known grooming incidents. These incidents were reported from real-world law enforcement case studies. The conversations in this dataset are labeled as predatory behaviors, coercion, and manipulative tactics. We use this dataset to train our context agent, accordingly. We also employ cyberbullying detection dataset [42, 43] which includes 4.5 million comments related to hate speech and cyberbullying. The dataset was collected from various social media platforms including Twitter, Reddit, and Discord. The dataset provides annotated comments for categories such as explicit content, threats, hate speech, profanity, and bullying. This dataset trains and assesses the sentinel agent’s detection accuracy. Before model training, the aforementioned datasets are preprocessed using tokenization, normalization, and filtering techniques. For testing purposes, we utilized GPT-4 to generate real-time simulated threat interactions with annotations to test how EdgeAIGuard adapts to new and evolving threat patterns that were not part of the training data. All the experiments are carried out on the simulated threat interactions from GPT-4.

### v2

@@ I Introduction
  • A Novel multi-agent framework EdgeAIGuard for minors’ online safety.
@@ II Related Work
The LLM-based approaches [33, 17, 12, 34, 11, 13] surpassing CNN and RNN models in detecting hate speech content in textual data. Authors in [33] explored the automatic detection of white supremacist hate speech on Twitter using the BERT architecture, which outperformed BiLSTM. Similarly, [17] introduced HateTinyLLM for efficient hate speech detection, using fine-tuned, decoder-only tiny large language models (tinyLLMs) combined with Low-Rank Adaptation (LoRA) and adapter methods. In [12], the authors presented SHIELD, an LLM-based technique that extracts rationales from input text to improve the interpretability of hate speech detection models. In [34], the researchers focused on enhancing hate speech detection with multilingual corpora and Cross-Lingual Learning (CLL) techniques, specifically for political discourse. Similarly, [11] investigated LLM-based methods for detecting sexist and hateful online content through information retrieval, employing zero-shot and few-shot learning approaches. Finally, [13] introduced TurkishBERTweet, a LLM-based method trained on over 894 million Turkish tweets. Recent work in cross-technology authentication may inform secure agent communication and edge-level verification, such as AUTHFi [35] and [36].
Current research methods on hate speech detection using machine learning, deep learning, and LLM-based methods have significant limitations when deployed on edge devices. Traditional models such as SVM [22] and feature-based methods [26, 28] struggle with context-dependent hate speech and require extensive feature engineering. Deep learning models like CNNs, RNNs, and LSTMs [31, 30, 8] improve classification compared to traditional approaches but suffer from high computational costs and limited generalization beyond their training datasets. While LLM-based methods [33, 11, 17] show promising results, they have significant latency issues, making real-time detection impractical. Due to high computational demands, existing LLM-based approaches are inefficient for edge deployment, raising concerns about privacy, cost, and real-time inference. Additionally, existing methods rely on specific datasets and fail to adapt to evolving online threats, limiting real-world applicability. We propose an Agentic AI-based approach designed on an edge-first architecture to address these challenges, ensuring enhanced privacy, minimal latency, and adaptive threat detection. The proposed EdgeAIGuard method is designed to detect grooming tactics such as trust building, isolation, desensitization, exclusivity, and manipulation via gifts or secrecy. It uses linguistic patterns and context history to identify such behaviors.
@@ III Methodology
The EdgeAIGuard framework comprises four essential layers, which are designed to optimize the protection pipeline for minors. The input layer collects data and processes information from various digital platforms. The form of information can be extended to the ones shown in Figure 2, such as comments on a blog, discord chat, or any other platform that can be used for communication with minors. The edge processing unit forms the analytical core of the EdgeAIGuard framework, housing both the agent network and LLM engine. Edge deployment significantly reduces privacy risk by keeping user data on-device entirely. Only severe threats require external reporting. Local storage and processing eliminate the need for cross-border data transfer, enhancing GDPR and COPPA compliance. Context-aware decision-making is performed through local storage, which maintains emerging patterns and historical data for the AI agents. Finally, the intervention in the case of threat detection is executed by the protection layer in synchronization with one of the AI agents.
The proposed system adopts a sequential-parallel hybrid processing approach by allowing multiple specialized agents to process incoming data and information from the communications simultaneously. The hybrid approach maintains critical sequential dependencies essential for decision-making. It also optimizes the accuracy and processing speed, thus making the proposed EdgeAIGuard compliant with the rapid response time requirement for threat detection systems. EdgeAIGuard includes fallback logic to preserve core functionality in case of agent unavailability. If the Sentinel agent fails, the Context agent uses pattern memory and stricter thresholds to maintain threat detection. If the Context agent is missing, single-message assessments are weighted more heavily with a temporary history buffer. If the Intervention agent fails, predefined responses based on threat severity are used. This redundancy ensures that minor protection remains active even with partial system degradation.
The intervention agent is designed to execute protective actions tailored to factors such as the probability of harmful content, incident history, and user vulnerability. The intervention agent ensures interventions are proportionate and effective, balancing protection with minimal disruption. The following are some examples of interventions generated at different severity levels while using the intervention agent.
III-A Low Severity
When the system detects mildly inappropriate language, the intervention agent may display a safety message, such as, ”The system detected some content that might not be appropriate. We advise being cautious and reporting anything that makes you uncomfortable.” Such intervention aims to alert the user and educate without disrupting the conversation, while logging the interaction for further monitoring.
III-B Moderate Severity
If potential grooming tactics are identified, such as emotionally manipulating or attempting to isolate the minor, the intervention agent might pause the chat temporarily while sending a notification to the platform moderator or guardian stating, ”We have detected behavior to isolate or manipulate the [minor’s name]. The interaction needs reviewing from your end, accordingly.” Such intervention will ensure the prompt involvement of guardians for additional oversight.
III-C High Severity
If a clear grooming attempt or explicit coercion is identified, the intervention agent will terminate the chat immediately and send an alert to the relevant authorities or the guardian with a message stating, ”Potential exploitation or grooming attempt targeting [minor’s name] detected. Your immediate attention is needed for the feedback. Details have been logged and encrypted for review.” Such a response will ensure that user safety is prioritized while providing a rapid response to critical risk.
The aforementioned examples inferred from the agent describe its functionality, integration, and protection explicitly. Furthermore, the agent’s ability to learn and adapt from feedback will enhance its accuracy in catering to evolving threats and users’ preferences.
III-D Preliminaries and Problem Formulation
@@ III-E Input Layer
III-F Edge Processing Unit
@@
The notation $H$ represents the historical context vector, $g$ is the scoring function, and the operator $\circ$ represents the fusion operation. Lastly, we define the intervention agent that determines protective actions based on threat severity and contextual risks, using a weighted assessment of harmful content probability, incident history, and user vulnerability. Interventions may include safety messages, chat pauses, or alerts to guardians, tailored to the user’s age and tone. The agent continually improves by evaluating response effectiveness and gathering feedback on false positives. We formally define the intervention agent in Equation 7.
@@ III-G Local Storage
The term $HC$ refers to the history cache, which utilizes a decay factor $\alpha$ and the current state represented by $CS(t)$, whereas the $pm_{z}$ refers to the set of patterns such that $(pattern,frequency,severity)$. The history cache maintains a record of recent system and communication states and is adjusted through time-weighting. The history cache also utilizes $\alpha$ to reduce the influence of older information gradually. However, it preserves important patterns accordingly. On the other hand, pattern memory stores recurring threat patterns, along with their severity and frequency. The memory enables the system to recognize evolving threat patterns and adapt responses. To protect privacy, all personal identifiers are anonymized using one-way hashing. We minimize and encrypt cached data using AES-256. Data processing is done entirely on-device, and all data is deleted after 30 days. The system complies with GDPR and COPPA by incorporating privacy-by-design principles, retaining data for only 7 days for minors under COPPA, and using features solely relevant to threats. The history cache retains anonymized interactions for 30 days, with older entries decaying gradually.
III-H Protection Layer
@@
The notations $TD$ refers to the threat decision, $risk$ corresponds to the historical risks, and $\delta$ refers to the decision function. The weights for each of the terms are optimized so that the decision process for intervention is optimal. Subsequently, the intervention selection process is designed to balance the effectiveness of potential interventions while analyzing the impact on the user experience. To mitigate false positives, EdgeAIGuard employs context-aware thresholds that adjust based on communication history and detected patterns. The Sentinel agent uses ensemble scoring from content, metadata, and interaction history to reduce single-signal misclassifications. The intervention agent incorporates user feedback, and frequently flagged false positives are logged to suppress similar future alerts via dynamic memory adjustments. In this regard, the intervention process employs a utility-based approach, as shown in Equation 13.
@@
III-I Objective Functions and Optimization
@@
where $loss_{ce}$ refers to the computational efficiency loss and $\lambda$s are the corresponding weights. The weights ($\lambda_{1}$$\lambda_{4}$) were determined via grid search followed by adaptive tuning. We define each loss component in Equations 15, 16, 17, and 18 as follows:
@@ IV-A Experimental Setup
The EdgeAIGuard was evaluated using a combination of software simulations, controlled testbeds, and benchmark datasets to replicate real-world scenarios. We used NVIDIA Jetson AGX Orin to emulate the edge device, representing real-world deployment conditions due to its balance between computational power and energy efficiency. EdgeAIGuard runs on Jetson AGX Orin device with a minimum specification of a 6-core ARM CPU, 8 GB of RAM, 384-core GPU, 32 GB of storage, and 10W of power. We also used Intel Xeon Gold 6226R CPU paired with an NVIDIA A100 GPU, representing a high-performance cloud environment for baseline comparisons. As mentioned in the algorithm and architectural workflow, we would leverage DeepSeek-7B (distilled version) [37], which is optimized for edge deployment without compromising performance. The DeepSeek-7B model was created using a teacher–student distillation framework. The 65B teacher model guided the student via output matching, intermediate representation alignment, and attention supervision. We employed INT8 quantization, attention pruning, and layer fusion. A distilled DeepSeek-7B model was used, achieving 87% size reduction and 73% latency reduction while preserving accuracy. For the model development and fine-tuning, we employed PyTorch [38] and HuggingFace Transformers [39].
@@ IV-B Dataset Details
We mainly used two datasets to train the sentinel and context agents to identify implicit grooming attempts and detect harmful interactions. We utilized the online grooming dataset [43] that contains 1.2 million chat messages from known grooming incidents. These incidents were reported from real-world law enforcement case studies. The conversations in this dataset are labeled as predatory behaviors, coercion, and manipulative tactics. We use this dataset to train our context agent, accordingly. We also employ cyberbullying detection dataset [44, 45] which includes 4.5 million comments related to hate speech and cyberbullying. The dataset was collected from various social media platforms including Twitter, Reddit, and Discord. The dataset provides annotated comments for categories such as explicit content, threats, hate speech, profanity, and bullying. This dataset trains and assesses the sentinel agent’s detection accuracy. Before model training, the aforementioned datasets are preprocessed using tokenization, normalization, and filtering techniques. For testing purposes, we utilized GPT-4 to generate real-time simulated threat interactions with annotations to test how EdgeAIGuard adapts to new and evolving threat patterns that were not part of the training data. All the experiments are carried out on the simulated threat interactions from GPT-4. To address dataset imbalance, we performed demographic audits and used GPT-4 to augment underrepresented groups. We applied stratified sampling across age, platform, and language categories. Adversarial training discouraged learning biased patterns.
@@ IV-C Experimental Analysis
EdgeAIGuard can support REST API by integrating existing moderation systems. It includes modules for platforms like Discord, Instagram, and TikTok to enable real-time alerts, webhook notifications, and family control extension. The proposed method is designed for scalable deployment across heterogeneous environments. The model architecture is modular and supports federated learning for decentralized updates. However, deployment at scale poses challenges such as variable hardware capabilities, update synchronization, and regulatory variations across jurisdictions. We mitigate these by providing device-specific model variants and centralized policy management interfaces for large platform partners. One of the limitations of the proposed work is the lack of evaluation on user-centered metrics such as adoption rates and user perceptions. As the evaluation of user-centered metrics is a large and diverse domain to cover, we would like to integrate the proposed work with a workable edge device in the future and provide it to users for real-world use. The aim would be to evaluate the user experience regarding trust, safety, and adoption. We intend to collect responses from minors and guardians on the system interventions. In addition, we would also like to track adoption rate, i.e., the number of users who enable the feature and retention over time, to understand the real-world impact.
@@ V Conclusion
One of the limitations of the proposed work is the lack of evaluation of user-centered metrics such as adoption rates and user perceptions. As the evaluation of user-centered metrics is a large and diverse domain to cover, we would like to integrate the proposed work with a workable edge device in the future and provide it to users for real-world use. The aim would be to evaluate the user experience regarding trust, safety, and adoption. We intend to collect responses from minors and guardians on the system interventions. In addition, we would also like to track the adoption rate, i.e., the number of users who enable the feature and retention over time, to understand the real-world impact.